# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

### v2

@@ Appendix B Length of Query Paths
$\inf_{\hat{f}}\max_{f\in\mathfrak{F}}\mathbb{E}\left\{\left\|\hat{f}-f\right\|_{2}^{2}\right\}\geq(\delta/2)^{2}\left(1-\frac{I(Z;J)+\log 2}{\log M}\right),$ (D.9)
@@
$\begin{split}\operatorname{KL}(p_{i}\|p_{j})&=\mathbb{E}\left\{\sum_{i=1}^{n}\operatorname{KL}\left(p_{i}(y_{i}|\mathbf{X}_{i})\|p_{j}(y_{i}|\mathbf{X}_{i})\right)\right\}\\ &=\mathbb{E}_{\mathbf{X}}\left\{\frac{n\left\|\delta\cdot f_{\mathbf{z}_{i}}-\delta\cdot f_{\mathbf{z}_{j}}\right\|_{n}^{2}}{2\sigma^{2}}\right\}\\ &=\frac{n\delta^{2}\left\|f_{\mathbf{z}_{i}}-f_{\mathbf{z}_{j}}\right\|_{2}^{2}}{2\sigma^{2}}\\ &\leq\frac{n\delta^{2}}{\sigma^{2}}.\end{split}$ (D.10)
@@
$I(Z;J)\leq\frac{1}{M^{2}}\sum_{i,j=1}^{M}\operatorname{KL}(p_{i}\|p_{j})\leq\frac{n\delta^{2}}{\sigma^{2}}.$ (D.11)
@@
$\frac{I(Z;J)+\log 2}{\log M}\leq\frac{1}{4}+\frac{\log 2}{2^{s-2}\log((d-s)/4)}\leq\frac{1}{2},$ (D.12)
@@
Lemma D.1.
Consider the set $[m]^{r}$ for some $m$ and $r$, equipped with the Hamming distance metric
@@
There is a $1/2$-separated set in $[m]^{r}$ of size at least $\exp\left(\frac{r}{2}\log(m/4)\right)$.
@@
$M(\delta;[m]^{r},d_{H})\geq\frac{\text{Vol}([m]^{r})}{\text{Vol}(B_{\delta})}$ (D.13)
where $B_{\delta}$ is a Hamming ball in $[m]^{r}$ of radius $\delta$. The volume of this ball satisfies
$\text{Vol}(B_{\delta})\leq\binom{r}{\delta r}m^{\delta r}.$ (D.14)
@@
$\begin{split}\log M(1/2;[m]^{r},d_{H})\geq r\log m-\frac{r}{2}\log m-r\log 2.\qed\end{split}$ (D.15)
Appendix E Proof of Proposition 6.2
@@
$\mathcal{G}_{n}(\delta;\mathfrak{F})\coloneqq\mathbb{E}_{\mathbf{W}}\left\{\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}f(\mathbf{X}_{i})\right|\right\},$ (E.1)
@@
¯$\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})\coloneqq\mathbb{E}_{\boldsymbol{\epsilon},\mathcal{D}_{n}}\left\{\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{2}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(\mathbf{X}_{i})\right|\right\}.$ (E.2)
We note that $\mathcal{G}_{n}(\delta;\mathfrak{F})$ is a random quantity (it still demands on $\mathcal{D}_{n}$) while ¯$\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})$ is fully deterministic. These quantities are used to provide bounds for nonparametric least squares estimation problems, which includes ERM over tree-based functions (11). It will also be helpful to define notation for the integrands in (E.1) and (E.2) as follows:
$\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})\coloneqq\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}f(\mathbf{X}_{i})\right|,$ (E.3)
¯$\bar{\mathcal{R}}_{n}(\boldsymbol{\epsilon};\delta;\mathfrak{F})\coloneqq\mathbb{E}_{\mathcal{D}_{n}}\left\{\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{2}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(\mathbf{X}_{i})\right|\right\}.$ (E.4)
Lemma E.1.
@@
$\mathfrak{F}\coloneqq\left\{\sum_{j=1}^{q}\alpha_{j}\phi_{j}\colon\boldsymbol{\alpha}\in\mathbb{R}^{q}\right\}.$ (E.5)
Then we have $\mathcal{G}_{n}(\delta;\mathfrak{F})\leq\delta(q/n)^{1/2}$ almost surely and ¯$\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})\leq\delta(q/n)^{1/2}$. Furthermore, we have the following tail bounds with respect to $\mathbf{W}$ and $\boldsymbol{\epsilon}$:
$\mathbb{P}\left\{\left|\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})-\mathcal{G}_{n}(\delta;\mathfrak{F})\right|\geq t\right\}\leq 2\exp\left(-\frac{nt^{2}}{2\delta^{2}}\right),$ (E.6)
{|¯$\mathbb{P}\left\{\left|\bar{\mathcal{R}}_{n}(\boldsymbol{\epsilon};\delta;\mathfrak{F})-\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})\right|\geq t\right\}\leq C_{1}\exp\left(-\frac{C_{2}nt^{2}}{2\delta}\right),$ (E.7)
@@
$\left\{\left(f(\mathbf{X}_{1}),f(\mathbf{X}_{2}),\ldots,f(\mathbf{X}_{n})\right)\colon f\in\mathfrak{F},\|f\|_{n}\leq\delta\right\}$ (E.8)
@@
$\mathcal{G}_{n}(\delta;\mathfrak{F})\leq\delta\sqrt{\frac{q}{n}}.$ (E.9)
@@
$\begin{split}\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{2}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(\mathbf{X}_{i})\right|&=\frac{1}{n}\sup_{\|\alpha\|_{2}\leq\delta}\left|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\boldsymbol{\alpha}\right|\\ &\frac{\delta}{n}\left\|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\right\|_{2}.\end{split}$ (E.10)
@@
¯$\begin{split}\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})&=\mathbb{E}_{\mathcal{D}_{n},\boldsymbol{\epsilon}}\left\{\frac{\delta}{n}\left\|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\right\|_{2}\right\}\\ &\leq\frac{\delta}{n}\mathbb{E}_{\mathcal{D}_{n},\boldsymbol{\epsilon}}\left\{\left\|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\right\|_{2}^{2}\right\}^{1/2}\\ &=\frac{\delta}{n}\mathbb{E}_{\mathcal{D}_{n}}\left\{\operatorname{Trace}\left(\boldsymbol{\Phi}^{T}\boldsymbol{\Phi}\right)\right\}^{1/2}\\ &=\delta\sqrt{\frac{q}{n}}.\end{split}$ (E.11)
@@
$\begin{split}\left|\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})-\mathcal{G}_{n}(\mathbf{W}^{\prime};\delta;\mathfrak{F})\right|&\leq\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\left(W_{i}-W_{i}^{\prime}\right)f(\mathbf{X}_{i})\right|\\ &\leq\frac{1}{n}\|\mathbf{W}-\mathbf{W}^{\prime}\|_{2}\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left\|f\right\|_{2}\\ &\leq\frac{\delta}{\sqrt{n}}\|\mathbf{W}-\mathbf{W}^{\prime}\|_{2},\end{split}$ (E.12)
where the second inequality follows from Cauchy-Schwarz. As such, the function $\mathbf{W}\mapsto\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})$ is $\delta/\sqrt{n}$-Lipschitz, and we may use Gaussian Lipschitz concentration (Wainwright, 2019) to obtain equation (E.6). Finally, we may use Talagrand’s convex concentration inequality (Talagrand, 1995) to get equation (E.7). ∎
@@
$\mathfrak{F}=\bigcup_{\mathcal{P}\in\mathfrak{P}_{s}}\mathfrak{F}_{\mathcal{P}}.$ (E.13)
@@
$\text{star}(\mathfrak{F}^{*},0)=\left\{\alpha(f-f^{*})\colon f\in\mathfrak{F},0\leq\alpha\leq 1\right\}.$ (E.14)
@@
$\text{star}(\mathfrak{F}^{*},0)=\bigcup_{\mathcal{P}\in\mathfrak{P}_{s}}\text{star}(\mathfrak{F}_{\mathcal{P}}^{*},0),$ (E.15)
observing that each term $\text{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)$ is a subset of a $12^{s}+$ 1-dimensional vector space. Moreover, we have
$\left|\mathfrak{P}_{s}\right|\leq d^{2^{s}},$ (E.16)
@@
$\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))\lesssim\delta\sqrt{\frac{2^{s}\log d}{n}},$ (E.17)
¯$\bar{\mathcal{R}}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))\lesssim\delta\sqrt{\frac{2^{s}\log d}{n}}.$ (E.18)
@@
$\begin{split}&\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))\\ &=\mathbb{E}_{\mathbf{W}}\left\{\sup_{\mathcal{P}\in\mathfrak{P}_{s}}\mathcal{G}_{n}\left(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)\right)\right\}\\ &\leq\mathbb{E}\left\{\left(\sum_{\mathcal{P}\in\mathfrak{P}_{s}}\mathcal{G}_{n}\left(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)\right)^{r}\right)^{1/r}\right\}\\ &\leq\left(\sum_{\mathcal{P}\in\mathfrak{P}_{s}}\mathbb{E}\left\{\mathcal{G}_{n}\left(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)\right)^{r}\right\}\right)^{1/r}\\ &\leq\left|\mathfrak{P}_{s}\right|^{1/r}\sup_{\mathcal{P}\in\mathfrak{P}_{s}}\left(\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))+\mathbb{E}\left\{\left|\mathcal{G}_{n}(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))-\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\right|^{r}\right\}^{1/r}\right).\end{split}$ (E.19)
Using equation E.6, we get
$\mathbb{E}\left\{\left|\mathcal{G}_{n}(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))-\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\right|^{r}\right\}^{1/r}\lesssim\delta\sqrt{\frac{r}{n}}.$ (E.20)
Combining this with the bound $\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\leq\delta\sqrt{\frac{2^{s}+1}{n}}$ from Lemma E.1 and selecting $r=2^{s}\log d$, we get
$\begin{split}\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))&\lesssim\delta\sqrt{\frac{2^{s+1}}{n}}+\delta\sqrt{\frac{2^{s}\log d}{n}}\\ &\lesssim\delta\sqrt{\frac{2^{s}\log d}{n}}.\end{split}$ (E.21)
Finally, equation (E.18) is proved similarly using (E.7) and the bound ¯$\bar{\mathcal{R}}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\leq\delta\sqrt{\frac{2^{s}+1}{n}}$. ∎
@@
$\delta_{n}\coloneqq\min\left\{\delta\colon\frac{\bar{\mathcal{R}}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))}{\delta}\leq\frac{\delta}{128C}\right\},$ (E.22)
$\epsilon_{n}\coloneqq\min\left\{\epsilon\colon\frac{\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))}{\epsilon}\leq\frac{\epsilon}{2\sigma}\right\},$ (E.23)
where $C$ is such that $\mathbb{E}\left\{f^{4}(\mathbf{x})\right\}\leq C^{2}\mathbb{E}\left\{f^{2}(\mathbf{x})\right\}$ for all $f\in\textnormal{star}(\mathfrak{F}^{*},0)$. Given our additional constraint on the supremum norm on functions in $\mathfrak{F}$, we can take $C=2M$. Since the additional constraint can only decrease the value of the local Gaussian and Rademacher complexities, Lemma E.2 gives upper bounds
$\epsilon_{n}\lesssim\sigma\sqrt{\frac{2^{s}\log d}{n}},\quad\delta_{n}\lesssim M\sqrt{\frac{2^{s}\log d}{n}}.$ (E.24)
@@
$R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\left(\sigma^{2}+M^{2}\right)\frac{2^{s}\log d}{n}$ (E.25)

### v3

@@ Appendix C Length of Query Paths
$\inf_{\hat{f}}\max_{f\in\mathfrak{F}}\mathbb{E}\left\{\left\|\hat{f}-f\right\|_{2}^{2}\right\}\geq(\delta/2)^{2}\left(1-\frac{I(Z;J)+\log 2}{\log M}\right),$ (E.9)
@@
$\begin{split}\operatorname{KL}(p_{i}\|p_{j})&=\mathbb{E}\left\{\sum_{i=1}^{n}\operatorname{KL}\left(p_{i}(y_{i}|\mathbf{X}_{i})\|p_{j}(y_{i}|\mathbf{X}_{i})\right)\right\}\\ &=\mathbb{E}_{\mathbf{X}}\left\{\frac{n\left\|\delta\cdot f_{\mathbf{z}_{i}}-\delta\cdot f_{\mathbf{z}_{j}}\right\|_{n}^{2}}{2\sigma^{2}}\right\}\\ &=\frac{n\delta^{2}\left\|f_{\mathbf{z}_{i}}-f_{\mathbf{z}_{j}}\right\|_{2}^{2}}{2\sigma^{2}}\\ &\leq\frac{n\delta^{2}}{\sigma^{2}}.\end{split}$ (E.10)
@@
$I(Z;J)\leq\frac{1}{M^{2}}\sum_{i,j=1}^{M}\operatorname{KL}(p_{i}\|p_{j})\leq\frac{n\delta^{2}}{\sigma^{2}}.$ (E.11)
@@
$\frac{I(Z;J)+\log 2}{\log M}\leq\frac{1}{4}+\frac{\log 2}{2^{s-2}\log((d-s)/4)}\leq\frac{1}{2},$ (E.12)
@@
Lemma E.1.
Consider the set $m]^{r}$ for some $m$ and $r$, equipped with the Hamming distance metric
@@
There is a $1/2$-separated set in $m]^{r}$ of size at least $\exp\left(\frac{r}{2}\log(m/4)\right)$.
@@
$M(\delta;[m]^{r},d_{H})\geq\frac{\text{Vol}([m]^{r})}{\text{Vol}(B_{\delta})}$ (E.13)
where $B_{\delta}$ is a Hamming ball in $m]^{r}$ of radius $\delta$. The volume of this ball satisfies
$\text{Vol}(B_{\delta})\leq\binom{r}{\delta r}m^{\delta r}.$ (E.14)
@@
$\begin{split}\log M(1/2;[m]^{r},d_{H})\geq r\log m-\frac{r}{2}\log m-r\log 2.\qed\end{split}$ (E.15)
Appendix F Proof of Proposition 6.2
@@
$\mathcal{G}_{n}(\delta;\mathfrak{F})\coloneqq\mathbb{E}_{\mathbf{W}}\left\{\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}f(\mathbf{X}_{i})\right|\right\},$ (F.1)
@@
$\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})\coloneqq\mathbb{E}_{\boldsymbol{\epsilon},\mathcal{D}_{n}}\left\{\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{2}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(\mathbf{X}_{i})\right|\right\}.$ (F.2)
We note that $\mathcal{G}_{n}(\delta;\mathfrak{F})$ is a random quantity (it still demands on $\mathcal{D}_{n}$) while $\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})$ is fully deterministic. These quantities are used to provide bounds for nonparametric least squares estimation problems, which includes ERM over tree-based functions (11). It will also be helpful to define notation for the integrands in (F.1) and (F.2) as follows:
$\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})\coloneqq\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}f(\mathbf{X}_{i})\right|,$ (F.3)
$\bar{\mathcal{R}}_{n}(\boldsymbol{\epsilon};\delta;\mathfrak{F})\coloneqq\mathbb{E}_{\mathcal{D}_{n}}\left\{\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{2}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(\mathbf{X}_{i})\right|\right\}.$ (F.4)
Lemma F.1.
@@
$\mathfrak{F}\coloneqq\left\{\sum_{j=1}^{q}\alpha_{j}\phi_{j}\colon\boldsymbol{\alpha}\in\mathbb{R}^{q}\right\}.$ (F.5)
Then we have $\mathcal{G}_{n}(\delta;\mathfrak{F})\leq\delta(q/n)^{1/2}$ almost surely and $\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})\leq\delta(q/n)^{1/2}$. Furthermore, we have the following tail bounds with respect to $\mathbf{W}$ and $\boldsymbol{\epsilon}$:
$\mathbb{P}\left\{\left|\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})-\mathcal{G}_{n}(\delta;\mathfrak{F})\right|\geq t\right\}\leq 2\exp\left(-\frac{nt^{2}}{2\delta^{2}}\right),$ (F.6)
$\mathbb{P}\left\{\left|\bar{\mathcal{R}}_{n}(\boldsymbol{\epsilon};\delta;\mathfrak{F})-\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})\right|\geq t\right\}\leq C_{1}\exp\left(-\frac{C_{2}nt^{2}}{2\delta}\right),$ (F.7)
@@
$\left\{\left(f(\mathbf{X}_{1}),f(\mathbf{X}_{2}),\ldots,f(\mathbf{X}_{n})\right)\colon f\in\mathfrak{F},\|f\|_{n}\leq\delta\right\}$ (F.8)
@@
$\mathcal{G}_{n}(\delta;\mathfrak{F})\leq\delta\sqrt{\frac{q}{n}}.$ (F.9)
@@
$\begin{split}\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{2}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(\mathbf{X}_{i})\right|&=\frac{1}{n}\sup_{\|\alpha\|_{2}\leq\delta}\left|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\boldsymbol{\alpha}\right|\\ &\frac{\delta}{n}\left\|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\right\|_{2}.\end{split}$ (F.10)
@@
$\begin{split}\bar{\mathcal{R}}_{n}(\delta;\mathfrak{F})&=\mathbb{E}_{\mathcal{D}_{n},\boldsymbol{\epsilon}}\left\{\frac{\delta}{n}\left\|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\right\|_{2}\right\}\\ &\leq\frac{\delta}{n}\mathbb{E}_{\mathcal{D}_{n},\boldsymbol{\epsilon}}\left\{\left\|\boldsymbol{\epsilon}^{T}\boldsymbol{\Phi}\right\|_{2}^{2}\right\}^{1/2}\\ &=\frac{\delta}{n}\mathbb{E}_{\mathcal{D}_{n}}\left\{\operatorname{Trace}\left(\boldsymbol{\Phi}^{T}\boldsymbol{\Phi}\right)\right\}^{1/2}\\ &=\delta\sqrt{\frac{q}{n}}.\end{split}$ (F.11)
@@
$\begin{split}\left|\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})-\mathcal{G}_{n}(\mathbf{W}^{\prime};\delta;\mathfrak{F})\right|&\leq\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\left(W_{i}-W_{i}^{\prime}\right)f(\mathbf{X}_{i})\right|\\ &\leq\frac{1}{n}\|\mathbf{W}-\mathbf{W}^{\prime}\|_{2}\sup_{\begin{subarray}{c}f\in\mathfrak{F}\\ \|f\|_{n}\leq\delta\end{subarray}}\left\|f\right\|_{2}\\ &\leq\frac{\delta}{\sqrt{n}}\|\mathbf{W}-\mathbf{W}^{\prime}\|_{2},\end{split}$ (F.12)
where the second inequality follows from Cauchy-Schwarz. As such, the function $\mathbf{W}\mapsto\mathcal{G}_{n}(\mathbf{W};\delta;\mathfrak{F})$ is $\delta/\sqrt{n}$-Lipschitz, and we may use Gaussian Lipschitz concentration (Wainwright, 2019) to obtain equation (F.6). Finally, we may use Talagrand’s convex concentration inequality (Talagrand, 1995) to get equation (F.7). ∎
@@
$\mathfrak{F}=\bigcup_{\mathcal{P}\in\mathfrak{P}_{s}}\mathfrak{F}_{\mathcal{P}}.$ (F.13)
@@
$\text{star}(\mathfrak{F}^{*},0)=\left\{\alpha(f-f^{*})\colon f\in\mathfrak{F},0\leq\alpha\leq 1\right\}.$ (F.14)
@@
$\text{star}(\mathfrak{F}^{*},0)=\bigcup_{\mathcal{P}\in\mathfrak{P}_{s}}\text{star}(\mathfrak{F}_{\mathcal{P}}^{*},0),$ (F.15)
observing that each term $\text{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)$ is a subset of a $2^{s}+1$-dimensional vector space. Moreover, we have
$\left|\mathfrak{P}_{s}\right|\leq d^{2^{s}},$ (F.16)
@@
$\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))\lesssim\delta\sqrt{\frac{2^{s}\log d}{n}},$ (F.17)
$\bar{\mathcal{R}}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))\lesssim\delta\sqrt{\frac{2^{s}\log d}{n}}.$ (F.18)
@@
$\begin{split}&\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))\\ &=\mathbb{E}_{\mathbf{W}}\left\{\sup_{\mathcal{P}\in\mathfrak{P}_{s}}\mathcal{G}_{n}\left(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)\right)\right\}\\ &\leq\mathbb{E}\left\{\left(\sum_{\mathcal{P}\in\mathfrak{P}_{s}}\mathcal{G}_{n}\left(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)\right)^{r}\right)^{1/r}\right\}\\ &\leq\left(\sum_{\mathcal{P}\in\mathfrak{P}_{s}}\mathbb{E}\left\{\mathcal{G}_{n}\left(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0)\right)^{r}\right\}\right)^{1/r}\\ &\leq\left|\mathfrak{P}_{s}\right|^{1/r}\sup_{\mathcal{P}\in\mathfrak{P}_{s}}\left(\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))+\mathbb{E}\left\{\left|\mathcal{G}_{n}(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))-\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\right|^{r}\right\}^{1/r}\right).\end{split}$ (F.19)
Using equation F.6, we get
$\mathbb{E}\left\{\left|\mathcal{G}_{n}(\mathbf{W};\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))-\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\right|^{r}\right\}^{1/r}\lesssim\delta\sqrt{\frac{r}{n}}.$ (F.20)
Combining this with the bound $\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\leq\delta\sqrt{\frac{2^{s}+1}{n}}$ from Lemma F.1 and selecting $r=2^{s}\log d$, we get
$\begin{split}\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))&\lesssim\delta\sqrt{\frac{2^{s+1}}{n}}+\delta\sqrt{\frac{2^{s}\log d}{n}}\\ &\lesssim\delta\sqrt{\frac{2^{s}\log d}{n}}.\end{split}$ (F.21)
Finally, equation (F.18) is proved similarly using (F.7) and the bound $\bar{\mathcal{R}}_{n}(\delta;\textnormal{star}(\mathfrak{F}_{\mathcal{P}}^{*},0))\leq\delta\sqrt{\frac{2^{s}+1}{n}}$. ∎
@@
$\delta_{n}\coloneqq\min\left\{\delta\colon\frac{\bar{\mathcal{R}}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))}{\delta}\leq\frac{\delta}{128C}\right\},$ (F.22)
$\epsilon_{n}\coloneqq\min\left\{\epsilon\colon\frac{\mathcal{G}_{n}(\delta;\textnormal{star}(\mathfrak{F}^{*},0))}{\epsilon}\leq\frac{\epsilon}{2\sigma}\right\},$ (F.23)
where $C$ is such that $\mathbb{E}\left\{f^{4}(\mathbf{x})\right\}\leq C^{2}\mathbb{E}\left\{f^{2}(\mathbf{x})\right\}$ for all $f\in\textnormal{star}(\mathfrak{F}^{*},0)$. Given our additional constraint on the supremum norm on functions in $\mathfrak{F}$, we can take $C=2M$. Since the additional constraint can only decrease the value of the local Gaussian and Rademacher complexities, Lemma F.2 gives upper bounds
$\epsilon_{n}\lesssim\sigma\sqrt{\frac{2^{s}\log d}{n}},\quad\delta_{n}\lesssim M\sqrt{\frac{2^{s}\log d}{n}}.$ (F.24)
@@
$R\left(\hat{f}_{\operatorname{ERM}}(-;\mathcal{D}_{n}),f^{*}\right)\leq\left(\sigma^{2}+M^{2}\right)\frac{2^{s}\log d}{n}$ (F.25)